OpenStudy (anonymous):

The speed of the current in a river is 6 mph. A ferry operator who works that part of the river is looking to buy a new boat for his business. Every day, his route takes him 22.5 miles against the current and back to his dock, and he needs to make this trip in a total of 9 hours. He has a boat in mind, but he can only test it on a lake where there is no current. How fast must the boat go on the lake in order for it to serve the ferry operator’s needs?

OpenStudy (anonymous):

@Nurali i need help on this

OpenStudy (merchandize):

When you travel 22.5 miles against the current, and then back the current opposes you one way and assists you the other way, so the effects cancel. So, the operators time on a lake with no current is the same as the time going up and down the river. --------------- Let s = the speed of the boat on the lake s=2*22.5/9 s=45/9 s=5 He needs to go 5 mi/hr on the lake for the boat to work on the river also

OpenStudy (anonymous):

I thought it would be 6mi/hr . I'm a little lost lol hold on i think im slowly getting it

OpenStudy (anonymous):

@merchandize ??????

OpenStudy (merchandize):

where you are not getting??

OpenStudy (anonymous):

is that what you got as a answer ?

OpenStudy (merchandize):

yes....5 is the answer

OpenStudy (anonymous):

Okay thank you. I just had a misunderstanding.

OpenStudy (merchandize):

okay... welcome...:)

OpenStudy (anonymous):

@SmokeysTheName right ?

OpenStudy (smokeysthename):

this is how i have mine

OpenStudy (anonymous):

Hmm okay let me make some changes lol

OpenStudy (anonymous):

I have that now. Whats up with 13

OpenStudy (smokeysthename):

ill tag you hold on

OpenStudy (anonymous):

okay